# Vacuum Energy and Topological Mass in Interacting Elko and Scalar Field Theories

### v1

@@ I Introduction
From the perspective of high-energy physics and cosmology, Elko fields have been extensively discussed as components of dark matter models, due to their suppressed couplings to ordinary matter and gauge bosons. They have also been considered in inflationary scenarios, as drivers of accelerated expansion, and in models of late-time cosmic acceleration where Elko condensates mimic dark energy. In curved spacetimes, Elko fields can modify gravitational dynamics through non-minimal couplings and play a role in ?????Olhar TEx code . These applications highlight their versatility and justify the study of Elko in diverse quantum and cosmological settings [27, 26, 24, 25]. In particular, Casimir-like effects associated with Elko fields offer a way to explore how exotic spinors interact with boundaries and topology [24, 25], thereby linking fundamental field-theoretical properties with potentially observable macroscopic phenomena.
@@ II Effective potential of the Elko field interacting with a real scalar field
where $\chi$ and $\rho$ represent quantum fluctuations, and $\Psi$ and $\Phi$ are fixed classical background fields. This procedure is analogous to what was done in Ref. [33]. In contrast, in the present case, due to the fermionic nature of the Elko field, there is no need to set $\Phi=0$ in advance, unlike in [33].
@@ III.2 Loop Correction and Generalized Zeta Functions
Here, the first term on the right-hand side of Eq. (33) is independent of the parameter $L$ and therefore does not contribute to the vacuum energy density. The second term increases with positive powers of the mass and should consequently be discarded [16]. Hence, only the third term contributes to the renormalized vacuum energy density, as we will see later.
@@
We can also, in principle, consider the case of a massless field only for the real scalar field. As a consequence of its nonlocal nature, the massless limit of the Elko field is singular (see Ref. [27]) and therefore not well-defined. An interesting additional feature of this field is that it becomes local as its mass increases. The massless limit of the real field has been investigated in several other works, for instance in Refs. [16, 31]. This limit can be obtained by considering, in Eq. (43),
@@ IV.1 First-Order Coupling-Constant Corrections to the Vacuum Energy
Note that the factor of three in Eq. (48) originates from Wickâaspastheorem for field contractions, while the symmetry factor in this case is $\tfrac{1}{2}$. The corresponding loop implies that the zeta function in Eq. (49) should appear divided by $\Omega_{4}$. Moreover, for the diagram under consideration there is a single vertex contributing with a factor $\tfrac{\lambda_{\varphi}}{4!}$.
@@ IV.2 Topological mass of the Elko field
Note that, due to the Grassmannian nature of the variables $\Psi\bar{\Psi}$, higher-order terms in Eq. (53) vanish.
@@ IV.3 Topological Mass of the Real Scalar Field
Note that one term of the correction is proportinal to the coupling constant $g$, since it comes from the interaction between the fields and it also depends on the mass of the Elko field. The other correction comes from the self-interaction of the real scalar field, so it is proportional to the coupling constant $\lambda_{\varphi}$. Moreover, the two terms also depends on the parameter $L$ related with the boundary condition.
@@ Acknowledgements.
The authors are grateful to Prof. D. Grumiller for useful discussions. H.F.S.M. is partially supported by the Brazilian agency National Council for Scientific and Technological Development (CNPq) under Grant No. 308049/2023-3. The author E.R.B.M. thanks CNPq for partial support, Grant No. 304332/2024-0.
@@ Appendix A On renormalization
After substituting the renormalization constants (63)–(65) into Eq. (62), the renormalized effective potential becomes
@@
where we take $(\bar{\Psi}\Psi)^{2}=0$ due to the fermionic nature of the Elko field, and

### v3

@@ I Introduction
From the perspective of high-energy physics and cosmology, Elko fields have been extensively discussed as components of dark matter models, due to their suppressed couplings to ordinary matter and gauge bosons. They have also been considered in inflationary scenarios, as drivers of accelerated expansion, and in models of late-time cosmic acceleration where Elko condensates mimic dark energy. These applications highlight their versatility and justify the study of Elko in diverse quantum and cosmological settings [27, 26, 24, 25]. In particular, Casimir-like effects associated with Elko fields offer a way to explore how exotic spinors interact with boundaries and topology [24, 25], thereby linking fundamental field-theoretical properties with potentially observable macroscopic phenomena.
@@ II Effective potential of the Elko field interacting with a real scalar field
where $\chi$ and $\rho$ represent quantum fluctuations, and $\Psi$ and $\Phi$ are fixed classical background fields. Note that $\Phi$ is a scalar field, while $\Psi$ is a single Grassmann variable, reflecting the fermionic nature of the Elko field. The procedure above is analogous to what was done in Ref. [33]. In contrast, in the present case, due to the fermionic nature of the Elko field, there is no need to set $\Phi=0$ in advance, unlike in [33].
@@ III.2 Loop Correction and Generalized Zeta Functions
Here, the first term on the right-hand side of Eq. (33) is independent of the parameter $L$ and therefore does not contribute to the vacuum energy density. The second term grows with positive powers of the mass and should therefore be discarded. This is consistent with a normalization condition requiring that the renormalized vacuum energy vanish in the large-mass limit [16]. Hence, only the third term contributes to the renormalized vacuum energy density, as we will see later.
@@
We may also, in principle, consider the massless limit only for the real scalar field. The fundamental anticommutation relations involving the Elko field operators depend on the inverse mass and are therefore ill-defined in the massless limit. This behavior is a consequence of the nonlocal nature of the Elko field (see Ref. [27]). An additional noteworthy feature of this field is that it becomes local as its mass increases.
The massless limit of the real field has been investigated in several other works, for instance in Refs. [16, 31]. This limit can be obtained by considering, in Eq. (43),
@@ IV.1 First-Order Coupling-Constant Corrections to the Vacuum Energy
Note that the factor of three in Eq. (48) originates from the Wick’s theorem for field contractions, while the symmetry factor in this case is $\tfrac{1}{2}$. The corresponding loop implies that the zeta function in Eq. (49) should appear divided by $\Omega_{4}$. Moreover, for the diagram under consideration there is a single vertex contributing with a factor $\tfrac{\lambda_{\varphi}}{4!}$.
@@ IV.2 Topological mass of the Elko field
Note that, due to the Grassmann nature of the variables $\Psi\bar{\Psi}$, which are proportional to a single Grassmann variable, all higher-order terms in Eq. (53) vanish.
@@ IV.3 Topological Mass of the Real Scalar Field
Note that one term of the correction is proportional to the coupling constant $g$, since it comes from the interaction between the fields and it also depends on the mass of the Elko field. The other correction comes from the self-interaction of the real scalar field, so it is proportional to the coupling constant $\lambda_{\varphi}$. Moreover, the two terms also depends on the parameter $L$ related with the boundary condition.
@@ Acknowledgements.
The authors are grateful to Prof. D. Grumiller and Prof. Andrea Erdas for useful discussions. H.F.S.M. is partially supported by the Brazilian agency National Council for Scientific and Technological Development (CNPq) under Grant No. 308049/2023-3. The author E.R.B.M. thanks CNPq for partial support, Grant No. 304332/2024-0.
@@ Appendix A On renormalization
After substituting the renormalization constants (63)-(65) into Eq. (62), the renormalized effective potential becomes
@@
where we take $(\bar{\Psi}\Psi)^{2}=0$, since it is proportional to a single Grassmann variable, and